# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

### v2

@@
Łukasz Debowski Łukasz Debowski is with the Institute of Computer Science, Polish Academy of Sciences, ul. Jana Kazimierza 5, 01-248 Warszawa, Poland, e-mail: ldebowsk@ipipan.waw.pl.
We construct multiperiodic processes—a simple example of stationary ergodic stochastic processes over natural numbers that enjoy the vanishing entropy rate under a mild condition. Multiperiodic processes are supported on randomly shifted deterministic sequences called multiperiodic sequences, which can be efficiently generated using an algorithm called the Infinite Clock. Under a suitable parameterization, multiperiodic sequences exhibit relative frequencies of particular numbers given by Zipf‘s law. Exactly in the same setting, the respective multiperiodic processes satisfy an asymptotic power-law growth of block entropy, called Hilberg‘s law. Hilberg‘s law is deemed to hold for statistical language models, in particular.
@@ 1 Introduction
The aim of this paper is to construct a simple stationary ergodic source with the vanishing entropy rate. In particular, we are interested in the block entropy obeying an asymptotic power law growth.
The notations are as follows. Let $H(X):=\operatorname{\mathbf{E}}\left[-\log P(X)\right]$ be the Shannon entropy of a discrete random variable $X$ where $\log$ is the base 2 logarithm, let $H(X|Y):=H(X,Y)-H(Y)$ be the conditional entropy, and let $I(X;Y):=H(X)-H(X|Y)$ be the mutual information. We assume a general familiarity with these Shannon information measures [1]. Relation $\sim$ denotes the equality of the asymptotic power law growth,
@@
Blocks of random variables are denoted $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$.
We are seeking for mathematical toy examples of stationary processes $(X_{t})_{t\in\mathbb{Z}}$ that obey
@@
where $\beta\in(0,1)$ and $h=\lim_{t\to\infty}H(X_{1}^{t})/t$ is the Shannon entropy rate. The rough proportionality (2) is called Hilberg‘s law after [2] and is deemed to hold for natural language [2, 3, 4, 5]. As it will be detailed in a separate paper, Hilberg‘s law implies the more famous neural scaling law, which is a power-law decay of the cross entropy rate of large language models [6, 7, 8].
Some simple known examples of processes that obey Hilberg‘s law (2) are Santa Fe processes [9, 4, 10]. The Santa Fe processes come in a number of varieties. Their common idea is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. In other words, each text token $X_{t}$ may be written as a pair
@@
The Santa Fe decomposition (3) is interesting for its abstract semantic interpretation, see [4, 5]. We may imagine that pairs $(K_{t},Z_{K_{t}})$ are logical statements that describe a certain unknown binary sequence $(Z_{k})_{k\in\mathbb{N}}$ at random but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may imagine that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable general knowledge which is only partially accessed and described in finite texts. For this reason, bits $Z_{k}$ appearing in representation (3) were called the ’’(elementary) facts‘‘, whereas sequence $(K_{t})_{t\in\mathbb{Z}}$ was called the ’’narration (process)‘‘ [5]. Thus, we may say that formula (3) decomposes a corpus of texts $(X_{t})_{t\in\mathbb{Z}}$ into ’’knowledge‘‘ $(Z_{k})_{k\in\mathbb{N}}$ and ’’narration‘‘ $(K_{t})_{t\in\mathbb{Z}}$.
@@
where $\alpha>1$ [11, 12]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (2) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^1] We obey the distinction of terms “type” and “token” borrowed from quantitative linguistics [13, 14, 15]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
@@
For the Zipf-Mandelbrot distribution (4), we have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha}$ [13, 16, 15]. Consequently, we may express
$H(X_{1}^{t})-H(K_{1}^{t})=H(\left\{(k,Z_{k}):k\in\mathcal{V}_{t}\right\}|K_{1}^{t})=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (6)
@@
The above described Santa Fe process is non-ergodic and takes values in a countably infinite alphabet. The reason for its non-ergodicity is that coin flips $(Z_{k})_{k\in\mathbb{N}}$ are measurable with respect to the shift-invariant algebra of process $(X_{t})_{t\in\mathbb{Z}}$. Thus each particular value of sequence $(Z_{k})_{k\in\mathbb{N}}$ defines a distinct ergodic component [9, 5]. However, replacing individual coin flips $Z_{k}$ with time-dependent but slowly evolving Markov chains $(Z_{kt})_{t\in\mathbb{Z}}$, we obtain processes which are (strongly) mixing hence ergodic [17]. Coding natural numbers as binary strings, we obtain processes over a finite alphabet [4, 18]. All these examples enjoy Hilberg‘s law (2) with a positive rate $h>0$.
@@
Our curiosity in such asymptotically deterministic models was initially driven by the observation that the maximal repetition length for natural language obeys a cube logarithmic growth [20]. Consequently, the conditional Rényi entropy rate may vanish for natural language [21]. Later, our interest was reinforced by the research in the neural scaling law [6, 7, 8], which suggested that Hilberg‘s law (2) holds for large language models with $h=0$. We are agnostic about the actual value of particular entropy rates for natural language but we have wished to better understand the mathematical construction of zero entropy rate processes.
@@
The essential idea of these processes is to define a probability measure on a set of efficiently computable but randomly shifted deterministic sequences, called multiperiodic sequences. Under a certain parameterization, the relative frequencies of particular numbers in these sequences approach the Zipf-Mandelbrot distribution (4). This suffices to imply Hilberg‘s law (7) with exponent $\beta=1/\alpha$ by a reasoning similar to bound (6). Namely, for the set of types (5), we still have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha}$ and we may express
$H(K_{1}^{t})=H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\sim\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (8)
where $(\Sigma_{k})_{k\in\mathbb{N}}$ are certain random variables called seeds. Seeds parameterize random shifts of the underlying deterministic multiperiodic sequences. Variables $(\Sigma_{k})_{k\in\mathbb{N}}$ are probabilistically independent but are not measurable with respect to the shift-invariant algebra of process $(K_{t})_{t\in\mathbb{Z}}$. Thus the multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ can be ergodic.
The organization of the paper is as follows. In Section 2, we define two objects: multiperiodic sequences in Section 2.1 and multiperiodic processes in Section 2.2. We analyze their statistics in Section 3: the relative frequencies in Section 3.1, the waiting times in Section 3.2, and the number of types in Section 3.3. In Section 4, we inspect the information flow: estimation of seeds given a finite sample in Section 4.1 and the block entropy in Section 4.2. These general developments are supplemented by examples in Section 5: constant periods in Section 5.1 and linear periods in Section 5.2. Section 6 concludes the article.
@@ 2.1 Multiperiodic sequences
Definition 1 (multiperiodic sequence)
@@
$\phi_{k,t+1}$ =$=\begin{cases}\phi_{k,t}-1,&k<k_{t},\\ \pi_{k},&k=k_{t},\\ \phi_{k,t},&k>k_{t},\end{cases}$ (11)
@@
$\phi_{k,t}$ =$=\begin{cases}\phi_{k,t+1}+1,&k<k_{t},\\ 1,&k=k_{t},\\ \phi_{k,t+1},&k>k_{t}.\end{cases}$ (13)
@@
Recursion (9)–(13) for $k\in\mathbb{N}$ can be rewritten as Algorithm 1, called the Infinite Clock. The Infinite Clock operates with infinite data structures for simplicity of presentation. It works like a clock with infinitely many hands—indexed by natural numbers—that move counterclockwise, hence the name. As in the usual clock, each hand moves with a different speed, dictated by its specific period. The initial positions of hands are given by seeds. The algorithm outputs the value of a hand index whenever the hand passes noon.
@@
  • Proof: Let $(\phi^{R}_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ and $(k^{R}_{t})_{t\in\mathbb{Z}}$ be defined analogously as (10)–(13) but with the initial condition $\phi^{R}_{k,1}=\sigma^{R}_{k}$. Then by induction we obtain $\phi^{R}_{k,t}=\pi_{k}-\phi_{k,1-t}+1$ and $k^{R}_{t}=k_{1-t}$. $\Box$
@@
  • Proof: We proceed by induction on $r$. As for the initial step, we have $(k^{\geq 1}_{t})_{t\in\mathbb{N}}=(k_{t})_{t\in\mathbb{N}}$. Thus claim 1. holds by definition, whereas claim 2. follows since $k_{t}=1$ if and only if $t\equiv\sigma_{1}\mod\pi_{1}$. As for the inductive step, it is sufficient to show that claim 1. holds, whereas claim 2. follows by claim 1. for the inductive step combined with claim 2. for the initial step. The proof of claim 1. for the inductive step rests on the fact that to obtain $(k^{\geq r}_{t})_{t\in\mathbb{N}}$ instead of $(k_{t})_{t\in\mathbb{N}}$, it suffices to initialize $k:=r$ instead of $k:=1$ in line 3 of Algorithm 1. $\Box$

### v3

@@
Łukasz Dębowski Łukasz Dębowski is with the Institute of Computer Science, Polish Academy of Sciences, ul. Jana Kazimierza 5, 01-248 Warszawa, Poland, e-mail: ldebowsk@ipipan.waw.pl.
We construct multiperiodic processes—a simple example of stationary ergodic (but not mixing) processes over natural numbers that enjoy the vanishing entropy rate under a mild condition. Multiperiodic processes are supported on randomly shifted deterministic sequences called multiperiodic sequences, which can be efficiently generated using an algorithm called the Infinite Clock. Under a suitable parameterization, multiperiodic sequences exhibit relative frequencies of particular numbers given by Zipf‘s law. Exactly in the same setting, the respective multiperiodic processes satisfy an asymptotic power-law growth of block entropy, called Hilberg‘s law. Hilberg‘s law is deemed to hold for statistical language models, in particular.
@@ 1 Introduction
The notations to state our problem are as follows. Blocks of random variables are denoted $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$. Let $H(X):=\operatorname{\mathbf{E}}\left[-\log P(X)\right]$ be the Shannon entropy of a discrete random variable $X$ where $\log$ is the base 2 logarithm, let H(X|Y):=H(X,Y)H(Y)H(X|Y):=H(X,Y)-H(Y) be the conditional entropy, and let I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y) be the mutual information [11]. Relation $\sim$ denotes is defined as
@@
Supplementing prior results, we seek for various toy examples of stationary processes $(X_{t})_{t\in\mathbb{Z}}$ that obey relationship
@@
where $\beta\in(0,1)$ is called the Hilberg exponent [19] and limit
$h:=\lim_{t\to\infty}\frac{H(X_{1}^{t})}{t}$ (3)
is the Shannon entropy rate. Relationship (2) is called Hilberg‘s law after [43]. The particular aim of the present paper is to construct simple examples of stationary ergodic (but not mixing) sources over a countable alphabet that obey Hilberg‘s law (2) with the vanishing entropy rate, $h=0$. The processes to be constructed are called multiperiodic processes and will be described after sketching the background.
To connect our problem to prior research, we note that limit
$E:=\lim_{t\to\infty}\left(H(X_{1}^{t})-ht\right)=\lim_{t\to\infty}I(X_{-t+1}^{0};X_{1}^{t})$ (4)
is called excess entropy [12] or predictive information [4, 3]. Excess entropy is an intuitive measure of memory stored in a stochastic process. Although this quantity only measures the memory capacity, without specifying how exactly the process future depends on the past, it can be given interesting interpretations. Some attempts were also made to generalize excess entropy to two-dimensional random fields [32, 8].
In fact, conditions $E<\infty$ and $E=\infty$ have been studied in probability theory and its applications for quite a time. The oldest systematic results seem to concern Gaussian processes. In particular, Grenander and Szegő [38, Section 5.5] present an integral formula for excess entropy (in disguise). Finch [33] evaluated this formula for autoregressive moving average (ARMA) processes, which turned out to yield $E<\infty$. More generally, under a positive and continuous spectral density, Li [49] showed that $E<\infty$ if and only if the autocorrelation $\rho_{k}:=\operatorname{Corr}(X_{0};X_{k})$ satisfies
$\sum_{k=1}^{\infty}k\rho_{k}^{2}<\infty.$ (5)
Dropping the condition for spectral density, we have $E<\infty$ if and only if $\sum_{k=1}^{\infty}k\alpha_{k}^{2}<\infty$ for the partial autocorrelation $\alpha_{k}$ [13].
Let us proceed to the countable alphabet. It can be easily shown, by the data-processing inequality, that condition $E<\infty$ holds for hidden Markov processes with a finite number of states [12]. Mahoney, Ellison and Crutchfield [52, 31] developed a formula for excess entropy in terms of predictive and retrodictive $\epsilon$-machines, which are minimal unifilar hidden Markov representations of the process [57, 51]. Travers and Crutchfield [58] and Dębowski [18] exhibited examples hidden Markov processes with a countably infinite number of hidden states that obey $E=\infty$ and (2). A few other older papers concerned also isolated examples of processes with $E=\infty$ over a finite alphabet. For instance, Bradley [7] constructed the first example of a mixing process having this property. Gramss [36] investigated a process which is formed by the frequencies of 0‘s and 1‘s in the rabbit sequence.
A more systematic study of condition $E=\infty$ for the countable alphabet case was inspired by the publication by Hilberg [43] suggesting that relationship (2) holds for natural language. As it will be detailed in a separate paper (to apply some results of the present one), Hilberg‘s law implies the more famous neural scaling law, which is a power-law decay of the cross entropy rate of large language models and empirically holds for internet-sized corpora of texts [47, 40, 42]. Although the first reception of Hilberg‘s hypothesis came from physicists [30, 29, 3, 12, 51], there is a line of research that seeks to combine a rigorous information-theoretic approach to Hilberg‘s law with general linguistic insights [14, 15, 16, 17, 24, 25] such as Zipf‘s law in particular [60, 53]. The basic observation is that condition $E=\infty$ holds if the process is strongly non-ergodic in the sense that its shift-invariant algebra $\mathcal{I}$ is non-atomic. It is so because
$E=H(\mathcal{I})+I(X_{-\infty}^{0};X_{1}^{\infty}|\mathcal{I})$ (6)
and $H(\mathcal{A})=\infty$ for a non-atomic algebra $\mathcal{A}$ under an appropriate generalization of Shannon information measures to arbitrary algebras of events [35, 28, 56, 59, 15, 23].
The ergodic decomposition of excess entropy (6) gives rise to a family of linguistically motivated examples of processes that obey Hilberg‘s law (2). These examples are called Santa Fe processes [15, 16] and were independently considered also by Hutter [44]. We recall this family of examples at length for their certain structural similarities and differences to the multiperiodic processes constructed in this paper. The Santa Fe processes come in a number of varieties—some are strongly non-ergodic, some are ergodic. Their common simple idea is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. In other words, each text token $X_{t}$ may be written as a pair
@@
The Santa Fe decomposition (7) is interesting for its abstract semantic interpretation, discussed in [16, 24]. We may interpret that pairs $(K_{t},Z_{K_{t}})$ are logical statements that describe a certain unknown binary sequence $(Z_{k})_{k\in\mathbb{N}}$ at random but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may interpret that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable general knowledge which is only partially accessed and described in finite texts. [^1] There are potential connections between the Santa Fe decomposition (7) and algorithmic information theory [50], see also [24, Chapter 1]. In the role of immutable knowledge, the random probability $Y:=\sum_{k=1}^{\infty}2^{-k}Z_{k}\in(0,1)$ conceptually resembles Chaitin’s number $\Omega\in(0,1)$, also called the halting probability [9, 10]. The number $\Omega$ is algorithmically random [54] and encodes a large body of mathematical knowledge in a dense form. However, equating the numerical values of probabilities $Y$ and $\Omega$ is preposterous for two reasons. The first of them, a formal one, is that $Y$ is a random variable, whereas $\Omega$ is a fixed number. The second reason, philosophically a more important one, is that $\Omega$ does not seem computable in a sufficiently strong sense to arise as an effect of natural processes [2]. For this reason, bits $Z_{k}$ appearing in representation (7) were called the ’’(elementary) facts‘‘ [16], whereas sequence $(K_{t})_{t\in\mathbb{Z}}$ was called the ’’narration (process)‘‘ [24]. Thus, we may say that formula (7) decomposes a corpus of texts $(X_{t})_{t\in\mathbb{Z}}$ into ’’knowledge‘‘ $(Z_{k})_{k\in\mathbb{N}}$ and ’’narration‘‘ $(K_{t})_{t\in\mathbb{Z}}$. To complete the naming, individual variables $K_{t}$ might be called ’’topics‘‘.
@@
where $\alpha>1$ [60, 53]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (2) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^2] We adopt the distinction of terms “type” and “token” borrowed from quantitative linguistics [41, 48, 1]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
@@
For the Zipf-Mandelbrot distribution (8), we have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha}$ [41, 39, 1]. Consequently, we may express
$H(X_{1}^{t})-H(K_{1}^{t})=H(\left\{(k,Z_{k}):k\in\mathcal{V}_{t}\right\}|K_{1}^{t})=\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha},$ (10)
@@
The above described Santa Fe process is non-ergodic and takes values in a countably infinite alphabet. The reason for its non-ergodicity is that coin flips $(Z_{k})_{k\in\mathbb{N}}$ are measurable with respect to the shift-invariant algebra of process $(X_{t})_{t\in\mathbb{Z}}$. Thus each particular value of sequence $(Z_{k})_{k\in\mathbb{N}}$ defines a distinct ergodic component [15, 24]. However, replacing individual coin flips $Z_{k}$ with slowly evolving Markov chains $(Z_{kt})_{t\in\mathbb{Z}}$, we obtain processes which are (strongly) mixing hence ergodic [17]. Coding natural numbers as binary strings, we obtain processes over a finite alphabet [16, 25]. All these examples enjoy Hilberg‘s law (2) with a positive rate $h>0$.
@@
Our curiosity in such asymptotically deterministic models was initially driven by the observation that the maximal repetition length for natural language obeys a cube logarithmic growth [20]. Consequently, the conditional Rényi entropy rate may vanish for natural language [22, 27]. Later, our interest was reinforced by the research in the neural scaling law [47, 40, 42], which suggested that Hilberg‘s law (2) holds for large language models with $h=0$. We are agnostic about the actual value of particular entropy rates for natural language but we have wished to better understand the mathematical construction of zero entropy rate processes.
@@
The essential idea of these processes is to define a probability measure on a set of efficiently computable but randomly shifted deterministic sequences, called multiperiodic sequences. Under a certain parameterization, the relative frequencies of particular numbers in these sequences approach the Zipf-Mandelbrot distribution (8). This suffices to imply Hilberg‘s law (11) with exponent $\beta=1/\alpha$ by a reasoning similar to bound (10). Namely, for the set of types (9), we still have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha}$ and we may express
$H(K_{1}^{t})=H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\sim\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha},$ (12)
where $(\Sigma_{k})_{k\in\mathbb{N}}$ are certain random variables called seeds. The seeds parameterize random shifts of the underlying deterministic multiperiodic sequences. Variables $(\Sigma_{k})_{k\in\mathbb{N}}$ are probabilistically independent and measurable with respect to process $(K_{t})_{t\in\mathbb{Z}}$ but are not measurable with respect to the shift-invariant algebra of process $(K_{t})_{t\in\mathbb{Z}}$. Thus the multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ can be ergodic. Moreover, we do not need to extend $(K_{t})_{t\in\mathbb{Z}}$ with additional coin flips $(Z_{k})_{k\in\mathbb{N}}$ and the Santa Fe construction (7) to obtain Hilberg‘s law. This remark concludes our brief presentation of the main results.
The organization of the paper is as follows. In Section 2, we define two objects: multiperiodic sequences in Section 2.1 and multiperiodic processes in Section 2.2. We analyze their statistics in Section 3: the relative frequencies in Section 3.1, the waiting times in Section 3.2, and the number of types in Section 3.3. In Section 4, we inspect the flow of information: estimation of seeds given a finite sample in Section 4.1 and the block entropy in Section 4.2. These general developments are supplemented by two examples in Section 5: constant periods in Section 5.1 and linear periods in Section 5.2. Section 6 concludes the article.
@@ 2.1 Multiperiodic sequences
Definition 1 (multiperiodic sequence).
@@
$\phi_{k,t+1}$ $=\begin{cases}\phi_{k,t}-1,&k<k_{t},\\ \pi_{k},&k=k_{t},\\ \phi_{k,t},&k>k_{t},\end{cases}$ (15)
@@
$\phi_{k,t}$ $=\begin{cases}\phi_{k,t+1}+1,&k<k_{t},\\ 1,&k=k_{t},\\ \phi_{k,t+1},&k>k_{t}.\end{cases}$ (17)
@@
Recursion (13)–(17) for $k\in\mathbb{N}$ can be rewritten as Algorithm 1, called the Infinite Clock. The Infinite Clock operates with infinite data structures for simplicity of presentation. It works like a clock with infinitely many hands—indexed by natural numbers—that move counterclockwise, hence the name. As in the usual clock, each hand moves with a different speed, dictated by its specific period. The initial positions of hands are given by seeds. The algorithm outputs the value of the hand index whenever a hand passes noon.
@@
Proof.
Let $(\phi^{R}_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ and $(k^{R}_{t})_{t\in\mathbb{Z}}$ be defined analogously as (14)–(17) but with the initial condition $\phi^{R}_{k,1}=\sigma^{R}_{k}$. Then by induction we obtain $\phi^{R}_{k,t}=\pi_{k}-\phi_{k,1-t}+1$ and $k^{R}_{t}=k_{1-t}$. ∎
@@
Proof.
We proceed by induction on $r$. As for the initial step, we have $(k^{\geq 1}_{t})_{t\in\mathbb{N}}=(k_{t})_{t\in\mathbb{N}}$. Thus claim 1. holds by definition, whereas claim 2. follows since $k_{t}=1$ if and only if $t\equiv\sigma_{1}\mod\pi_{1}$. As for the inductive step, it is sufficient to show that claim 1. holds, whereas claim 2. follows by claim 1. for the inductive step combined with claim 2. for the initial step. The proof of claim 1. for the inductive step rests on the fact that to obtain $(k^{\geq r}_{t})_{t\in\mathbb{N}}$ instead of $(k_{t})_{t\in\mathbb{N}}$, it suffices to initialize $k:=r$ instead of $k:=1$ in line 3 of Algorithm 1. ∎